OpenStudy (anonymous):

Find the speeds of two trains that are 80 miles apart if one train is traveling 10 mph faster than the other train and they pass each other in 25 min.

OpenStudy (anonymous):

slower train = 182 mph faster train = 192 mph tell me if it is wrong

OpenStudy (anonymous):

x and x+10 (2x+10)*25/60 = 80 => 2x + 10 = 80*60/25 => x = 91 so, 91 and 101 mph....

OpenStudy (anonymous):

ummmm not sure

OpenStudy (anonymous):

i m sure, my solution is right.....

OpenStudy (anonymous):

yep wrote my equation wrong thanks so much

OpenStudy (anonymous):

u welcome....

OpenStudy (anonymous):

i just got confused the direction of the train.. the train come in the opposite direction right. and then sudhansu is right :)